OpenStudy (anonymous):

factoring trinomials?! 10m^2-13mn-3n^2

Directrix (directrix):

@kayladavidson Do you know how to factor 10m^2 - 13m - 3 ? If so, we'll begin there.

Directrix (directrix):

Wow, I was about to respond and your answer vaporized.

OpenStudy (anonymous):

Well I knew it wasn't right. I've been doing problems like this all night nothing seems to make sense anymore.

Directrix (directrix):

Let's step back. Does factoring x^2 - 5x + 6 seem possible for you to do? You would look for two numbers that multiply to 6 and add to -5.

OpenStudy (anonymous):

-6 and 1. I get things like that, but when it's something like the one I'm trying now, I don't understand what can add up to -13 and multiply to -3

Directrix (directrix):

I hear you. First, -6 and 1 do not multiply to 6. They do add to -5. The numbers you want are -3 and -2. Then, factor by grouping. x^2 - 5x + 6 = x^2 - 3x -2x +6 = x(x-3) - 2(x -3) = (x-3)(x-2) That's the kind of thing we're about to do with the problem you posted.

Directrix (directrix):

10m^2-13mn-3n^2 ----------------- Needed: two numbers that multiply to -30 and add to -13 ARE the following: -15 and +2. Agree?

Directrix (directrix):

10m^2-13mn-3n^2 = 10m^2 -15mn + 2 mn - 3n^2 = ? Do you see how to do the next step?

OpenStudy (anonymous):

Factor by grouping?

Directrix (directrix):

Yes.

OpenStudy (anonymous):

5m(2m-3n)n(2m-3n) (5m+n)(2m-3n) ?

Directrix (directrix):

10m^2-13mn-3n^2 = 10m^2 -15mn + 2 mn - 3n^2 = 5m( 2m - 3n) + n(2m -3n) = The line above is the grouping. The object is to get the same factor in each expression. In this problem, (2m - 3n) is part of both expressions and will factor out.

Directrix (directrix):

Will you pick up from here: 5m( 2m - 3n) + n(2m -3n) = Think about it as factoring a common factor from: 2x + 4y which would give 2(x + 2y). That may or may not help.

OpenStudy (anonymous):

(5m+n)(2m-3n)

Directrix (directrix):

5m( 2m - 3n) + n(2m -3n) = (2m - 3n) (5m + n) which is what you have.

OpenStudy (anonymous):

Thanks!

Directrix (directrix):

Glad to help. It takes time and practice to get the hang of factoring. I recall getting extremely frustrated when I first started to learn how to do it. So, don't give up.